OpenStudy (anonymous):

Show by use of an example, that: (note the strict inequality)\[\lim_{n \rightarrow \infty} \sup (x _{n}+y _{n}) < \lim_{n \rightarrow \infty} \sup (x _{n})+\lim_{n \rightarrow \infty} \sup (y _{n}) \] and also \[\lim_{n \rightarrow \infty} \inf (x _{n}+y _{n}) > \lim_{n \rightarrow \infty} \inf (x _{n})+\lim_{n \rightarrow \infty} \inf (y _{n}) \]

OpenStudy (anonymous):

Where \[x _{n}\ and \y _{n}\] are bounded sequences.

OpenStudy (anonymous):

Take \[ x_n=(-1)^n\\ y_n=(-1)^{n+1}\\ x_n+y_n=0\\ \lim \sup x_n =1\\ \lim \sup y_n =1\\ \lim \sup (x_n+ y_n)=0\\ 0 < 2\\ \lim \inf x_n =-1\\ \lim \inf y_n =-1\\ \lim \inf (x_n+ y_n)=0\\ -2 <0 \]